OpenStudy (anonymous):

how to find the slope of this equation x + y = 7

OpenStudy (anonymous):

ax+by+c=0 , the slope is -a/b so answer is -1

OpenStudy (anonymous):

ok, were did the other variables "a" and "c" and "by" come from?? im completely confused now. where did the 7 go? and how did you get -1??

OpenStudy (anonymous):

'a' and 'c' not variables. x+y-7=0 then you should do -x's coefficient / y's coefficient , so -1/1=-1

OpenStudy (anonymous):

y=mx+b where m is slope. So x+y=7 y=-x+7=(-1)x + 7 m=-1 :D

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

cleared up some confusion